Find the solution fort in the equation 5(21)=4. O t= -0.322 Ot=0.861 O t= -0.861 Ot=0.322
Answer #1

5(2)t= 4

2t= 4/5

Taking log both sides

t log 2= log(4/5)

t= log (0.8)/log (2)

t= - 0.322